OpenStudy (anonymous):

HELP!!! Circle P has a radius of 8 units with center P at (2, -1). Which equation defines circle P? (x + 2)2 + (y – 1)2 = 64 (x – 2)2 + (y + 1)2 = 64 (x – 2)2 + (y + 1)2 = 8 (x + 2)2 + (y – 1)2 = 8

OpenStudy (anonymous):

circle formula \[(x-h)^2+(y-k)^2=r^2\] center of circle is (h,k) with radius r

OpenStudy (anonymous):

??

OpenStudy (anonymous):

that is the formula for a circle..sorry i'm not quite sure how else to explain it

OpenStudy (anonymous):

its ight ill just get someone els to help me buh thnks thoe [; !

OpenStudy (raden):

i just continue what cherio said, (x-h)^2 + (y-k)^2 = r^2 center of circle is (h,k) with radius r here given the center is (2,-1) and its radius = 8. So, just put the values of them : (x-2)^2 + (y-(-1))^2 = 8^2 (x-2)^2 + (y+1)^2 = 64